# Contextual AD Narration with Interleaved Multimodal Sequence

### v1

@@ Abstract
The Audio Description (AD) task aims to generate descriptions of visual elements for visually impaired individuals to help them access long-form video contents, like movie. With video feature, text, character bank and context information as inputs, the generated ADs are able to correspond to the characters by name and provide reasonable, contextual descriptions to help audience understand the storyline of movie. To achieve this goal, we propose to leverage pre-trained foundation models through a simple and unified framework to generate ADs with interleaved multimodal sequence as input, termed as Uni-AD. To enhance the alignment of features across various modalities with finer granularity, we introduce a simple and lightweight module that maps video features into the textual feature space. Moreover, we also propose a character-refinement module to provide more precise information by identifying the main characters who play more significant role in the video context. With these unique designs, we further incorporate contextual information and a contrastive loss into our architecture to generate more smooth and contextual ADs. Experiments on the MAD-eval dataset show that Uni-AD can achieve state-of-the-art performance on AD generation, which demonstrates the effectiveness of our approach. Code will be available at https://github.com/MCG-NJU/Uni-AD.
Figure 1: Task illustration. Taking video clip, text, character bank and context information as the inputs, the narrator generates corresponding audio description (AD) for video comprehension. Rather than describe all characters appearing in the video, the narrator should focus on characters that truly contribute to the storyline.
$*{}^{*}$ Intern at Ant Group. $\dagger$ Corresponding author.
@@ 1 Introduction
Audio Description (AD) [1, 13, 39] provides descriptive narration of visual content in videos. Unlike subtitle or transcription, AD focuses more on describing the scene, characters, actions and storyline of the input video. As a rich visual description, AD can effectively supplement the dialogue and provide viewers with a comprehensive description of the video content, which not only helps the visual impairments better engage with video content[1], but also benefits the sighted individuals in their media comprehension activities[34, 20], such as language learning for kids and sight-free video consuming while driving. Despite that AD is important for video comprehension, particularly for those professionally produced media contents (movies, documentary films, TV series etc.), currently most videos do not have corresponding AD[6], mainly due to the considerable costs of manual annotation and differences in understanding between annotators[14]. Therefore, studying how to generate ADs automatically is quite meaningful and necessary.
@@
With the advances in computer vision and natural language processing, nowadays the research community is paying growing attention to generating ADs automatically, which requires a model to be able to understand multi-modal information and perform contextual reasoning over video storyline[15, 46]. Compared with the conventional video captioning task[18, 10, 11, 29, 37], Audio Description (AD) is not only a scene description of the video clip, but also a narration that includes characters’ names and actions to generate a coherent plot description. This brings two characteristics of the AD generating task: (i) Multiple modality inputs. Video clip, text, character portraits and names are provided for AD generating. (ii) Rich contextual information. Context video and past AD can be applied to assist the current AD generation. An example is shown in Fig. 1.
@@
Previous methods[15, 14] introduce learnable adapters into GPT-2 to generate visual-conditioned ADs. However, the amount of parameters in these adapters will rapidly increase with larger LLM[14], thus not conducive to scaling up. MM-Narrator[46] directly prompts GPT-4 with specialized expert tools for AD generation, which suffers from complex prompt engineering and hallucination problem. Instead, in this work, we propose Uni-AD, a simple architecture that takes interleaved multimodal sequence as input to leverage completely open-source LLMs[36, 43] for AD generation by aligning various modality inputs to a unified semantic space. Such a structure brings multiple benefits. First, formulating data as interleaved multimodal sequence makes it convenient to integrate various modality inputs and add contextual information for AD generation. Second, Uni-AD converts visual inputs into the unified feature space with a lightweight module, thus we can efficiently take more powerful LLMs as decoder to generate more detailed and reasonable ADs. Third, interleaved multimodal sequence converts visual elements into multiple embedding tokens while maintaining the relative order between data, which ensures embeddings and tokens with the same semantic are naturally close to each other, so that more fine-grained feature alignment can be learned spontaneously.
@@
Given that the development of storyline is always character-centered[19, 42], it is necessary for ADs to include character names to describe their expressions, actions and status. Previous method [14] tried to identify all characters appearing in the given video as character information, without considering who are the main story drivers that should be included in AD. For example, in Fig. 1, though there are multiple characters in the scene, the expression change of Lisa is the main content thus only her name is involved in the target AD. With such an observation, we in this paper design a character-refinement module to figure out the AD-related characters. After training, this module can be applied to any videos to recognize main characters who contribute to the storyline and provide more accurate character information.
@@
To sum up, we make the following contributions:
  • By formulating multiple inputs as interleaved multimodal sequence, we develop an AD narration system called Uni-AD which achieves finer-grained feature alignment and supports extension to larger LLMs.
  • We introduce a character-refinement module and incorporate contextual information along with a contrastive loss to produce more accurate, coherent audio descriptions.
  • Our model significantly outperforms previous methods on the MAD-eval dataset. With the lightweight mapping network, we first extend our method to the fully available billion-scale LLaMA2-7B[43] and the experiment result shows further improvement, demonstrating the scalability of our approach.
@@ 2.1 Audio Descriptions Generation
Audio Description (AD) describes the key visual elements in videos to form coherent storyline narration. With the development of media technology, captioning for videos has emerged as a growing area in the computer vision research community[2, 25, 18]. Nonetheless, the production of Audio Description (AD) for video content remains a relatively untapped area of research. Initial works designed specialized authoring tools[7] and evaluation mechanisms[26, 30] to collect manually annotated ADs. Several annotation platforms like LiveDescribe[7], Rescribe[33] also emerged to facilitate AD generation. Since these manual methods require high labeling costs, they can’t scale well to the vast amount of visual media. Recently, some works have studied how to generate AD at scale automatically with deep learning models. AutoAD-I[15] introduced the task of AD generation for movies and addressed it by prompting GPT-2 with learnable visual prompt vectors. Researchers further incorporated an external character bank[14] to provide character information for more accurate AD generation. Zhang et al. [46] proposed a training-free design which extracts information from inputs with multimodal experts and queries GPT-4[31] in a few-shot manner. However, these methods suffer from drawbacks like poor scalability, complex prompt engineering and weak modality alignment, which we in this paper address with the interleaved multimodal sequence design and larger LLM.
@@ 2.3 LLMs for Video Understanding
The recent surge in large language models (LLMs)[47, 43, 36, 8, 31] has inspired the study of video perception and understanding with LLMs. Models like ChatCaptioner[9], VideoChat[22] and MM-Vid[24] integrate visual experts with LLMs to construct multimodal perception systems for video representation, long-term video comprehension and dialog-centric interaction, etc. All these works can be divided into three categories: (i) Prompt tuning[28, 9] is a lightweight approach to transfer LLMs to downstream tasks with learnable prompt vectors. (ii) Adapter-based methods[22] typically insert additional trainable parameters into the LLM at different positions to achieve deep modality alignment, but the amount of introduced parameters increases with the size of LLM, making adapter-based methods difficult to scale up. (iii) Querying LLMs in a training free manner[24], which employs visual experts to transform video into text, thereby guiding LLMs in reasoning on specific tasks. However, the effectiveness of this method diminishes when there is a significant gap between the downstream task and pre-training data. AD generation has a certain gap with training data of LLMs, hence the training-free methods are not very applicable. We think that more powerful pre-trained LLM leads to better description, thus we follow the visual-conditioned prompt tuning manner to extend our approach to larger LLMs with its memory-friendly characteristic.
@@ 3 Methodology
The audio description (AD) task is challenging mainly due the requirement that generated results should include characters’ names to depict their expressions, actions to advance the plot and should be contextually coherent within the storyline. To meet this, we present a unified model, Uni-AD, to formulate various inputs for current video clip along with contextual information as interleaved multimodal sequence and combines a character-refinement module. In this section, we will give a detailed description of our approach. First, we provide an overview of Uni-AD in Sec. 3.1. Then we present how to prompt a LLM for AD generation in Sec. 3.2. Next, we show the design and training details of our proposed character-refinement module in Sec. 3.3. Afterwards, we describe the usage of contextual information for AD generation in Sec. 3.4. Finally, in Sec. 3.5 we further introduce the learning object of our framework.
@@ 3.1 Method Overview
3.2 AD Generation with Interleaved Multimodal Prompt
3.2.1 Visual Mapping Network.
Figure 3: Structure of visual mapping network. MHSA denotes Multi-Head Self-Attention layer and FFN denotes FeedFoward layer.
The visual input for AD generation includes two components: (1) video clip consisting of $N$ frames with timestamp $t$, denoted as $v_{t}=\{\mathcal{I}_{1},\mathcal{I}_{2},...,\mathcal{I}_{N}\}$; (2) characters’ portrait images, denoted as $\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{C}\}$, where $C$ is the number of related characters in video clip $v_{t}$. To produce corresponding AD for $v_{t}$, we need to transform visual elements into embeddings with visual mapping network to achieve cross-modal alignment. Inspired by ClipCap[28] and AutoAD-I[15], we here apply a multi-layer transformer encoder with a fixed number of learnable vectors as our mapping network based on the following findings: First, video frames are able to interact with each other to model the temporal relation via the attention mechanism, which is essential given that effective interaction between visual tokens is hard to achieve in a frozen LLM. Second, the introduction of learnable vectors allows us to control the length of the visual representation and using more embeddings to preserve visual details. Finally, such a structure requires no addition of adapters inserted into the LLM for modality alignment, making it easier to scale to larger LLMs compared to methods like AutoAD-II[15]. The structure of our visual mapping network is shown in Fig. 3.
@@ 3.2 AD Generation with Interleaved Multimodal Prompt
$z_{v_{t}}=f_{CLIP}(v_{t}),\quad\{z_{1},z_{2},...,z_{C}\}=f_{CLIP}(\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{C}\}).$ (1)
@@
$x_{v_{t}}=\mathcal{M}_{v}(Proj(z_{v_{t}})),$ (2)
$\{x_{1},x_{2},...,x_{C}\}=\mathcal{M}_{v}(Proj(e_{1}),Proj(e_{2}),...,Proj(e_{C})),$ (3)
where $e_{i}$ denotes the $i_{th}$ character’s exemplar feature and $Proj$ represents a Linear Layer that transforms the channel number of visual features to match the LLM. Note that exemplar features are mapped via $\mathcal{M}_{v}$ separately, meaning there is no need for interaction between characters in this phase.
3.2.2 Formulating Interleaved Multimodal Prompt.
With visual embeddings $x_{v_{t}}$ and $\{x_{1},x_{2},...,x_{C}\}$, we now combine them with the text query to get our interleaved multimodal prompt. We apply the prompting template in [14] to query the frozen LLM, and our prompt is formulated as:
<Possible characters: $char_{1}$ played by $actor_{1}$ $x_{1}$, $char_{2}$ played by $actor_{2}$ $x_{2}$, … Describe $x_{v_{t}}$:>,
@@ 3.3 Character-Refinement Module
Figure 4: Structure of character-refinement module. MHCA denotes Multi-Head Cross-Attention layer and FFN denotes FeedFoward layer.
@@ 3.4 Contextual Information Modeling
Context ADs contain descriptions of preceding story plot that leads up to the current scene, thus can help the model better follow the storyline for narration. We here utilize context AD by directly concatenating our prompt in Sec. 3.2.2 with the past $K$ ADs $\{\mathcal{T}_{t-K},\mathcal{T}_{t-K},...,\mathcal{T}_{t-1}\}$. In this way, we provide more text conditional information for AD generation. To separate context ADs from current AD, we add a $BOS$ token in our prompt to start AD generation.
@@
$x_{t\_context}=\mathcal{M}_{v}(Proj(z_{v_{t-K}};z_{v_{t-K+1}};...;z_{v_{t-1}};z_{v_{t}})),$ (4)
@@
$s=\frac{\sum_{n}logP_{\Theta}(a_{n}|\mathrm{prompt};a_{<n})}{||\mathcal{T}||},\quad\mathcal{L}_{ct}=max(0,s_{last}-s_{current}),$ (5)
where $\mathcal{T}$ denotes an AD, $a_{n}$ denotes the $n_{th}$ token in $\mathcal{T}$, $a_{<n}$ denotes tokens preceding $a_{n}$ in $\mathcal{T}$; prompt denotes our interleaved multimodal sequence fed to LLM; [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] denotes concatenation; $\Theta$ denotes learnable parameters in our visual mapping network; $s$ represents the average likelihood score of the generated AD. Our contrastive loss $\mathcal{L}_{ct}$ is calculated as Eq. 5. Thus our model is constrained to ensure that the score of generating last AD ($s_{last}$) is always lower than score of the current ground-truth AD ($s_{current}$). In this way, we encourage the model to generate more accurate and non-repetitive ADs.
@@ 4 Experiments
In this section, we evaluate Uni-AD on the AD generation benchmark established in AutoAD[15] and show the experiment results. We first introduce our implementation details in Sec. 4.1. Then we compare our performance with state-of-the-art AD generation approaches in Sec. 4.2. We further conduct a detailed ablation study on the impact of character-refinement module, visual mapping network and partial-pre-training on our model in Sec. 4.3. Next in Sec. 4.4, we confirm the effectiveness of incorporating contextual information into Uni-AD and show how different contextual information affects our model. Finally, we provide qualitative examples of our Uni-AD in Sec. 4.5.

### v2

@@ Abstract
The Audio Description (AD) task aims to generate descriptions of visual elements for visually impaired individuals to help them access long-form video content, like movies. With video feature, text, character bank and context information as inputs, the generated ADs are able to correspond to the characters by name and provide reasonable, contextual descriptions to help audience understand the storyline of movie. To achieve this goal, we propose to leverage pre-trained foundation models through a simple and unified framework to generate ADs with interleaved multimodal sequence as input, termed as Uni-AD. To enhance the alignment of features across various modalities with finer granularity, we introduce a simple and lightweight module that maps video features into the textual feature space. Moreover, we also propose a character-refinement module to provide more precise information by identifying the main characters who play more significant roles in the video context. With these unique designs, we further incorporate contextual information and a contrastive loss into our architecture to generate smoother and more contextually appropriate ADs. Experiments on multiple AD datasets show that Uni-AD performs well on AD generation, which demonstrates the effectiveness of our approach. Our code is available at: https://github.com/ant-research/UniAD.
footnotetext: Corresponding author.
@@ 1 Introduction
Audio Description (AD) [1, 14, 40, 48] provides descriptive narration of visual content in videos. Unlike subtitle or transcription, AD focuses more on describing the scene, characters, actions and storyline of the input video. As a rich visual description, AD can effectively supplement the dialogue and provide viewers with a comprehensive description of the video content, which not only helps the visual impairments better engage with video content [1], but also benefits the sighted individuals in their media comprehension activities [35, 21], such as language learning for kids and sight-free video consuming while driving. Despite that AD is important for video comprehension, particularly for those professionally produced media contents (movies, TV series etc.), currently most videos do not have corresponding AD [5], mainly due to the considerable costs of manual annotation and differences in understanding between annotators [16]. Therefore, studying how to generate ADs automatically is quite meaningful and necessary.
@@
With the advances in computer vision and natural language processing, nowadays the research community is paying growing attention to generating ADs automatically, which requires a model to understand multi-modal information and perform contextual reasoning over video storyline [15, 49]. Compared with the conventional video captioning task [19, 9, 11, 30, 38], Audio Description (AD) is not only a scene description of the video clip, but also a narration that includes characters’ names and actions to generate a coherent plot description, as shown in LABEL:fig:teaser. This brings two characteristics of the AD generating task: (i) Multiple modality inputs. Video clip, text, character portraits and names are provided for AD generating. (ii) Rich contextual information. Context video and past AD can be applied to assist the current AD generation.
@@
Previous methods [15, 16] introduce learnable adapters into GPT-2 to generate ADs. However, the amount of parameters in these adapters will rapidly increase with larger LLM [16], thus not conducive to scaling up. Instead, AutoAD-III [17] applied Q-former architecture to bridge the visual space with the language space. Training-free methods like MM-Narrator [49] and AutoAD-Zero [48] directly prompt GPT-4 or LLaMA3 [13] with specialized expert tools for AD generation, which suffers from complex prompt engineering and hallucination problem. In this work, we propose Uni-AD, a simple architecture that takes interleaved multimodal sequence as input to leverage completely open-source LLMs [36, 44] for AD generation by aligning various modality inputs to a unified semantic space. Formulating data as interleaved multimodal sequence makes it convenient to integrate various modality inputs and add contextual information for AD generation. Besides, interleaved multimodal sequence converts visual elements into multiple embedding tokens while maintaining the relative order between data, which ensures embeddings and tokens with the same semantic are naturally close to each other so that more fine-grained feature alignment can be learned spontaneously.
@@
Given that the development of storyline is always character-centered [20, 43], it is necessary for ADs to include character names to describe their expressions, actions and status. Previous method [16] tried to identify all characters appearing in the given video as character information, without considering who are the main story drivers that should be included in AD. For example, in LABEL:fig:teaser, though there are multiple characters in the scene, the expression change of Lisa is the main content thus only her name is involved in the target AD. With such an observation, we in this paper design a character-refinement module to figure out the AD-related characters. After training, this module can be applied to any videos to recognize main characters who contribute to the storyline and provide more accurate character information.
@@
To sum up, we develop an AD narration system called Uni-AD which achieves finer-grained feature alignment and supports extension to larger LLMs by formulating multiple inputs as interleaved multimodal sequence. To produce more accurate, coherent audio descriptions, we introduce a character-refinement module and incorporate contextual information along with a contrastive loss. Our Uni-AD outperforms previous methods on multiple AD datasets.
@@ 2.1 Audio Descriptions Generation
Audio Description (AD) describes the key visual elements in videos to form coherent storyline narration. With the development of media technology, captioning for videos has emerged as a growing area in the computer vision research community [2, 25, 19]. Nonetheless, the production of Audio Description (AD) for video content remains a relatively untapped area of research. Initial works designed specialized authoring tools [6] and evaluation mechanisms [27, 31] to collect manually annotated ADs. Several annotation platforms like LiveDescribe [6], Rescribe [34] also emerged to facilitate AD generation. Recently, some works have studied how to generate AD at scale automatically with deep learning models. AutoAD-I [15] introduced the task of AD generation for movies and addressed it by prompting GPT-2 with learnable visual prompt vectors.  [16] later incorporated an external character bank to provide character information for more accurate AD generation. Researchers further applied Q-former architecture to bridge the visual and language space [17] for this task. Training free methods [49, 48] proposed designs which extract information from inputs with multimodal experts and queries GPT-4 [32] or VideoLLaMA2 [10] in a few-shot manner. However, these methods suffer from drawbacks like poor scalability, complex prompt engineering and weak modality alignment, which we in this paper address with the interleaved multimodal sequence design and larger LLM.
@@ 2.3 LLMs for Video Understanding
The recent surge in large language models (LLMs) [51, 44, 36, 7, 32] has inspired the study of video perception and understanding with LLMs. Models like ChatCaptioner [8], VideoChat [23] and MM-Vid [26] integrate visual experts with LLMs to construct multimodal perception systems for video representation, long-term video comprehension and dialog-centric interaction, etc. All these works can be divided into three categories: (i) Prompt tuning [29, 8] is a lightweight approach to transfer LLMs to downstream tasks with learnable prompt vectors. (ii) Adapter-based methods [23] typically insert additional trainable parameters into the LLM at different positions to achieve deep modality alignment, but the amount of introduced parameters increases with the size of LLM, making adapter-based methods difficult to scale up. (iii) Querying LLMs in a training free manner [26], which employs visual experts to transform video into text, thereby guiding LLMs in reasoning on specific tasks. Our Uni-AD follows the visual-conditioned prompt tuning manner to extend our approach to larger LLMs with its memory-friendly characteristic to generate better ADs.
@@ 3 Methodology
The audio description (AD) task is challenging mainly due the requirement that generated results should include characters’ names to depict their expressions, actions to advance the plot and should be contextually coherent within the storyline. To meet this, we present Uni-AD to formulate various inputs for current video clip along with contextual information as interleaved multimodal sequence and combines a character-refinement module. In this section, we will give a detailed description of our approach. First, we provide an overview of Uni-AD in Sec. 3.1. Then we present how to prompt a LLM for AD generation in Sec. 3.2. Next, we show the design and training details of our proposed character-refinement module in Sec. 3.3. Afterwards, we describe the usage of contextual information for AD generation in Sec. 3.4. Finally, in Sec. 3.5 we further introduce the learning object of our framework.
@@ 3.1 Method Overview
3.2 Uni-AD Pipeline
Visual Mapping Network. The visual input for AD generation includes two components: (1) video clip consisting of $N$ frames with timestamp $t$, denoted as $v_{t}=\{\mathcal{I}_{1},\mathcal{I}_{2},...,\mathcal{I}_{N}\}$; (2) characters’ portrait images, denoted as $\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{C}\}$, where $C$ is the number of related characters in video clip $v_{t}$. To produce corresponding AD for $v_{t}$, we need to transform visual elements into embeddings with visual mapping network to achieve cross-modal alignment. Inspired by ClipCap [29] and AutoAD-I [15], we here apply a multi-layer transformer encoder with a fixed number of learnable vectors as our mapping network (shown in Fig. 2) based on the following findings: First, video frames are able to interact with each other to model the temporal relation via the attention mechanism, which is essential given that effective interaction between visual tokens is hard to achieve in a frozen LLM. Second, the introduction of learnable vectors allows us to control the length of the visual representation and using more embeddings to preserve visual details. Finally, such a structure requires no addition of adapters inserted into the LLM for modality alignment, making it easier to scale to larger LLMs compared to methods like AutoAD-II [15].
@@ 3.2 Uni-AD Pipeline
$\begin{split}z_{v_{t}}&=f_{CLIP}(v_{t}),\\ \{z_{1},z_{2},...,z_{C}\}&=f_{CLIP}(\{\mathcal{A}_{1},\mathcal{A}_{2},...,\mathcal{A}_{C}\}).\end{split}$ (1)
@@
$\begin{split}x_{v_{t}}&=\mathcal{M}_{v}(Proj(z_{v_{t}})),\\ \{x_{1},x_{2},...,x_{C}\}&=\mathcal{M}_{v}(Proj(e_{1}),...,Proj(e_{C})),\end{split}$ (2)
where $e_{i}$ denotes the $i_{th}$ character’s exemplar feature and $Proj$ represents a Linear Layer that transforms the channel number of visual features to match the LLM. Note that exemplar features are mapped via $\mathcal{M}_{v}$ separately, meaning there is no need for interaction between characters here.
Formulating Interleaved Multimodal Prompt. With visual embeddings $x_{v_{t}}$ and $\{x_{1},x_{2},...,x_{C}\}$, we now combine them with the text query to get our interleaved multimodal prompt. We apply the prompting template in  [16] to query the frozen LLM, and our prompt is formulated as:
$\langle$ Possible characters: $char_{1}$ played by $actor_{1}$ $x_{1}$, $char_{2}$ played by $actor_{2}$ $x_{2}$, … Describe $x_{v_{t}}$: $\rangle$,
@@
Figure 2: Structure of Visual Mapping Network and Character-Refinement Module.
@@ 3.4 Contextual Information Modeling
Context ADs contain descriptions of preceding story plot that leads up to the current scene, thus can help the model better follow the storyline for narration. We here utilize context AD by directly concatenating our prompt in Sec. 3.2 with the past $K$ ADs $\{\mathcal{T}_{t-K},\mathcal{T}_{t-K},...,\mathcal{T}_{t-1}\}$. In this way, we provide more text conditional information for AD generation. To separate context ADs from current AD, we add a $BOS$ token in our prompt to start AD generation.
@@
$x_{t\_context}=\mathcal{M}_{v}(Proj(z_{v_{t-K}};z_{v_{t-K+1}};...;z_{v_{t}})),$ (3)
@@
$\begin{split}s=\frac{\sum_{n}logP_{\Theta}(a_{n}|\mathrm{prompt};a_{<n})}{||\mathcal{T}||},\\ \mathcal{L}_{ct}=max(0,s_{last}-s_{current}),\end{split}$ (4)
where $\mathcal{T}$ denotes an AD, $a_{n}$ denotes the $n_{th}$ token in $\mathcal{T}$, $a_{<n}$ denotes tokens preceding $a_{n}$ in $\mathcal{T}$; prompt denotes our interleaved multimodal sequence fed to LLM; $\Theta$ denotes learnable parameters in our visual mapping network; $s$ represents the average likelihood score of the generated AD. Our contrastive loss $\mathcal{L}_{ct}$ is calculated as Eq. 4. Thus our model is constrained to ensure that the score of generating last AD ($s_{last}$) is always lower than score of the current ground-truth AD ($s_{current}$). In this way, we encourage the model to generate more accurate and non-repetitive ADs.
@@ 4 Experiments
In this section, we evaluate Uni-AD on multiple AD generation benchmarks and show the experiment results. We first introduce our implementation details in Sec. 4.1. Then we compare our performance with state-of-the-art AD generation approaches in Sec. 4.2. We further conduct a detailed ablation study on the impact of character-refinement module and visual mapping network on our model in Sec. 4.3. Next in Sec. 4.4, we confirm the effectiveness of incorporating contextual information into Uni-AD and show how different contextual information affects our model. Finally, we provide qualitative examples of our Uni-AD in Sec. 4.5.